OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (anonymous):

First find g(2). Then find f of that number.

OpenStudy (anonymous):

g(2) = 2^2 - 6(2) - 7 = 4 - 12 - 7 = - 15

OpenStudy (anonymous):

Now, you find f(-15).

OpenStudy (anonymous):

f(-15) means to find the value of the function f when x = -15.

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

f(-15)=-15+8 f(-15)=23

OpenStudy (ranga):

You have to watch the sign of a number. 15 + 8 = 23 But -15 + 8 is NOT 23. Try again.

OpenStudy (anonymous):

-7

OpenStudy (ranga):

Yes.